OpenStudy (anonymous):

What are the approximate solutions of 2x2 − 7x = 3 rounded to the nearest hundredth? No real solutions x ≈ −0.77 and x ≈ 7.77 x ≈ −0.39 and x ≈ 3.89 x ≈ −3.89 and x ≈ 0.39

OpenStudy (campbell_st):

rewrite the equation \[2x^2 - 7x - 3 = 0\] use the general quadratic formula \[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] in your question a = 2, b = -7 and c = -3 evaluate to 3 decimal places for your answer

OpenStudy (anonymous):

what do you mean by "evaluate to 3 decimal places for your answer"?

OpenStudy (precal):

three places after the decimal sign

OpenStudy (anonymous):

so i plug in thhose numbers into that equation?? and then when i have the answer i bring it 3 decimal places to the right?

OpenStudy (campbell_st):

substitute the numbers and then get you answers to 3 decimal places

OpenStudy (campbell_st):

thats it remember there are 2 answers.....

OpenStudy (precal):

http://www.purplemath.com/modules/rounding.htm should be 2 places for hundredth

OpenStudy (campbell_st):

I think the 3rd option will be close to the answer... just doing some mental arithmetic

OpenStudy (campbell_st):

oops.. you only need 2 decimal places.... I just re-read the question

OpenStudy (anonymous):

i did it..but i came out with a weird equation....

OpenStudy (campbell_st):

do you have radicals in your answer..?

OpenStudy (campbell_st):

you need to find the value of \[x = \frac{7 + \sqrt{73}}{4}... and ... x = \frac{7 - \sqrt{73}}{4}\]

OpenStudy (anonymous):

7+- squared 25 ---------------- 4

OpenStudy (anonymous):

wait how did you get 73??

OpenStudy (campbell_st):

ok... a couple of things you need to put the -7 in brackets before squaring \[\sqrt{(-7)^2 - 4 \times2\times-3} = \sqrt{49 + 24} = \sqrt{73}\]

OpenStudy (campbell_st):

calcuators see the negative as an operator.... and not as part of the number

OpenStudy (anonymous):

-4 times 2 would be 8? then 8 times -3 would get what though?

OpenStudy (campbell_st):

-4 x 2 = -8 -8 x -3 = 24

OpenStudy (anonymous):

yeah so itsnt it 24 not 73?

OpenStudy (campbell_st):

. thats it

OpenStudy (anonymous):

oh pellet nvm!! i see you add the 49

OpenStudy (campbell_st):

yes... -7 x -7 = 49

OpenStudy (anonymous):

so then what do we do after that?

OpenStudy (campbell_st):

ok so you have \[x = \frac{7 + \sqrt{73}}{4}\] get a value for this correct to 2 decimal places then get a value for \[x = \frac{7 - \sqrt{73}}{4}\]

OpenStudy (anonymous):

how do i even put that into a decimal?

OpenStudy (campbell_st):

in your calculator just type the numerator and press equals... then its answer divided by 4 if you have a graphing calculator you may need 2nd = to get the decimal approximation the answers are 3.89 and 0.39

OpenStudy (anonymous):

thank you!!

OpenStudy (anonymous):

but wait there has to be one negative cause thats not an answer...

OpenStudy (anonymous):

so is it C?

OpenStudy (precal):

3.886 -.386

OpenStudy (precal):

all the work above is correct, just dropped sign on .386

OpenStudy (campbell_st):

oops yes... I missed a negative

OpenStudy (anonymous):

alright so it is answer C? thats what you guys are saying...

OpenStudy (campbell_st):

yes thats what we are saying